Assume the average price for a movie is $8.21. Assume the population standard deviation is $0.58 and that a sample of 34 theaThe answers are in red! Thank you in advance
Answer #1

u= 8.21, 0 = 0.58 , n = 34

a) Standard Error :

σε -

0.58 oT= = 0.0995 V34

b)PT <8.38)

Z = 2

8.38 – 8.21 Z = 0.099951.7091

Plī <8.38) = P(Z < 1.7091) = 0.9563

#value of z is obtain from standard normal table

c)Plī < 8.15)

Z = 2

= 8.15-8.21 0.09995 -=-0.6032 一

PI< 8.15) = P(Z < -0.6032) = 0.2732

d)PT> 8.33)

Z = 2

Z = 8.33 - 8.21 0,09995 -= 1.2064

PT> 8.33) = P(Z > 1.2064) = 1- P(Z < 1.2064)

  =1-0.8862 = 0.1138

PT> 8.33) = 0.1138